OpenStudy (anonymous):

the first term of an arithmetic progression a_1,a_2,a_3,....equal to unity. At what value of difference of the progression is a_1a_3+a_2a_3 is minimum

OpenStudy (shubhamsrg):

let a2 =a then a1 = a-d a3 = a+d we are given a-d =1 we need to minimize (a-d)(a+d) + a(a+d) easier now sire?

OpenStudy (anonymous):

thanks a lot..

OpenStudy (shubhamsrg):

Glad I could help.